1-3 Displacement 17 A traveler flies from Spokane to Atlanta by way of os Angeles. Los Angeles is approximately 1600 km due south of Spokane, and Atlanta is approximately 3200 km due east of Los Angeles (Fig. 1-32). Find the travelers total displacement NADA Spok NITED STA Los Angeles Atlanta MEXICO Scale 500 1000 miles 05001000 kilometers Fig. 1-32
Answer #1

displacement = minimum distance between starting point and ending point

In this case the direct line connecting Spokane and Atlanta

So given that

d1 = distance between Spokane and LA = 1600 km due south = (-1600 km) j

d2 = distance between LA and Atlanta = 3200 km due east = (3200 km) i

So distance will be given by vector addition of d1 and d2

d = d1 + d2

d = -1600 j + 3200 i

displacement will be

|d| = sqrt (d1^2 + d2^2)

|d| = sqrt ((-1600)^2 + 3200^2)

|d| = 3577.7 km

Please Upvote.